determine a possible quadratic function for points (2,7), (3,5) and (5,13).
Found 2 solutions by josgarithmetic, Alan3354:
Answer by josgarithmetic(39617)

You have a choice.
x=ay^2+by+c if you want symmetry axis parallel to the x-axis;
y=ax^2+bx+c if you want symmetry axis parallel to the y-axis;
or something with xy term if some other symmetry axis.

Each given point gives a separate, specific equation. Make THREE such equations, and solve for a, b, c, in the linear system. Do you need to see this done or started?

Answer by Alan3354(69443)

determine a possible quadratic function for points (2,7), (3,5) and (5,13).
-----------
2 parabolas can be found that pass thru the points.
---
y = ax^2 + bx + c is one.
---
4a + 2b + c = 7
9a + 3b + c = 5
25a + 5b + c = 13
----------
Eqn2 - Eqn 1:
5a + b = -2 **** Eqn 4
----------------
Eqn 3 - Eqn 2:
16a + 2b = 8
10a + 2b = -4 Eqn 4 times 2
---------------------------------- Subtract
6a = 12
a = 2
============
b = -12
=============
c = 23
================
f(x) = 2x^2 - 12x + 23
=======================================================
Find the 2nd one the same way, y as the independent variable.